# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 2 Main Results
To begin, we compute the gradient of the objective function as follows:
$\nabla_{\bm{W}_{l}}F(\bm{W})=\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\bm{Y}\right)\bm{W}_{l-1:1}^{T}+\lambda_{l}\bm{W}_{l},\ \forall l\in[L].$
Then, the set of critical points of Problem (2) is defined as
$\mathcal{W}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla_{\bm{W}_{l}}F(\bm{W})=\bm{0},\ \forall l\in[L]\right\}.$ (4)
Before we proceed, it is important to highlight the challenges in characterizing the set of critical points and establishing the error bound inequality (3) for Problem (2). First, the objective function is non-convex, presenting a complex loss landscape that may include a variety of critical points such as global minimizers, local minimizers, and saddle points [31]. Second, the objective function admits a hierarchical structure, where the weight matrices are multiplied sequentially and the layers have varying sizes. This hierarchical composition introduces a rotational invariance to the solution set (4), i.e., $\bm{W}_{l+1}\bm{W}_{l}=(\bm{W}_{l+1}\bm{Q}^{T})(\bm{Q}\bm{W}_{l})$ for any $\bm{Q}\in\mathcal{O}^{d_{l}}$ while the objective remains unchanged. This invariance leads to equivalence classes of solutions, thereby significantly complicating the analysis.
2.1 Characterization of Critical Points
In this subsection, we characterize the critical point set of Problem (2). To proceed, we first introduce the following assumption on the width of network layers.
@@ 2.1 Characterization of Critical Points
It holds that $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$.
This assumption ensures that the width of each hidden layer is no less than that of the input or output layer. It is important to note that this assumption is not restrictive and aligns with common practices in deep learning, where hidden layers are often designed to have sufficient capacity to capture complex data representations [15]. In addition, this assumption is widely used in the literature to analyze the convergence behavior of GD for optimizing deep networks; see, e.g., [3, 10, 11, 27].
Throughout the rest of the paper, let $d_{\min}:=\min\{d_{0},d_{1},\dots,d_{L}\}$. This, together with 1, implies $d_{\min}=\min\{d_{0},d_{L}\}$. Let $r_{Y}:=\mathrm{rank}(\bm{Y})$ and
$\bm{Y}=\bm{U}_{Y}\bm{\Sigma}_{Y}\bm{V}_{Y}^{T}$ (5)
be a singular value decomposition (SVD) of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$, where $\bm{U}_{Y}\in\mathcal{O}^{d_{L}}$, $\bm{V}_{Y}\in\mathcal{O}^{d_{0}}$, and $\bm{\Sigma}_{Y}=\mathrm{BlkDiag}(\mathrm{diag}(y_{1},y_{2},\dots,y_{r_{Y}}),\bm{0})\in\mathbb{R}^{d_{L}\times d_{0}}$ with $y_{1}\geq y_{2}\geq\dots\geq y_{r_{Y}}>0$ being top $r_{Y}$ positive singular values. In the literature, it is common to assume that the singular values of $\bm{Y}$ are distinct to simplify the analysis; see, e.g., [1, 24]. However, this assumption is overly strict and does not hold in many practical scenarios. For example, when $\bm{Y}$ is a membership matrix in $K$-classification problems, it typically has $K$ repeated singular values. To address this challenge, we introduce the following key setup and notions to improve the analysis. Let $p_{Y}$ denote the number of distinct positive singular values of $\bm{Y}$. In other words, there exist indices $s_{0},s_{1},\dots,s_{p_{Y}}$ such that $0=s_{0}<s_{1}<\dots<s_{p_{Y}}=r_{Y}$ and
$y_{s_{0}+1}=\dots=y_{s_{1}}>y_{s_{1}+1}=\dots=y_{s_{2}}>\dots>y_{s_{p_{Y}-1}+1}=\dots=y_{s_{p_{Y}}}>0,$ (6a)
$y_{s_{p_{Y}}+1}=\cdots=y_{d_{\text{min}}}=0.$ (6b)
Then, let $h_{i}:=s_{i}-s_{i-1}$ denote the multiplicity of the $i$-th largest positive singular value for each $i\in[p_{Y}]$. Consequently, we have $\sum_{i=1}^{p_{Y}}h_{i}=r_{Y}$. Now, we are ready to characterize the set of critical points explicitly as follows:
Theorem 1.
Suppose that 1 holds. It holds that $\bm{W}\in\mathcal{W}$ if and only if
$\bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right)\bm{V}_{Y}^{T},$ (7a)
@@
$\bm{W}_{L}=\bm{U}_{Y}\mathrm{BlkDiag}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},$ (7c)
where $\bm{Q}_{l}\in\mathcal{O}^{d_{l-1}}$ for all $l=2,\dots,L$, $\bm{O}_{i}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$, $\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}}$, $\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}$, and $\bm{\Sigma}_{l}=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma})/\sqrt{\lambda_{l}},\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for each $l\in[L]$ with $\bm{\sigma}\in\mathbb{R}^{r_{Y}}$ satisfying
$\sigma_{i}^{2L-1}-\left(\sqrt{\lambda_{1}\dots\lambda_{L}}y_{i}\right)\sigma_{i}^{L-1}+\left(\lambda_{1}\dots\lambda_{L}\right)\sigma_{i}=0,\ \sigma_{i}\geq 0,\ \forall i\in[r_{Y}].$ (8)
Before we proceed, let us make some remarks about this theorem. First, despite the non-convexity and hierarchical structure of Problem (2), we still characterize the set of critical points of Problem (2) explicitly, where each weight matrix admits the SVD in (7). Note that all weight matrices share the same singular values defined in (8) up to scaling of the regularization parameters $\{\lambda_{l}\}_{l=1}^{L}$, and their left and right singular matrices are determined by the orthogonal matrices $\bm{Q}_{l}$ and $\bm{O}_{i}$. Here, $\bm{Q}_{l}$ for all $l$ are introduced to handle the rotational invariance in the sequential matrix multiplication of the weight matrices, while $\bm{O}_{i}$ for all $i$ are used to address the repeated singular values in $\bm{Y}$.
Second, to the best of our knowledge, there is no complete characterization of the critical point set for regularized deep linear networks in the existing literature. Recently, Dang et al. [7] have studied the geometric properties of the global minimizers of the neural collapse problem with deep linear networks. Notably, this problem is a special instance of Problem (2), where $\bm{Y}$ is a membership matrix. In contrast to their work, which focuses only on global minimizers, our approach provides an explicit characterization of all critical points for arbitrary $\bm{Y}$. In addition, considerable research has been conducted to study the critical points of unregularized deep linear networks. The most complete and recent result is [1], which provides both necessary and sufficient conditions for identifying first-order critical points. Nevertheless, their analysis is limited to unregularized settings and cannot be applied to our case due to regularization terms.
Third, the explicit characterization of the critical point set in (7) serves as a cornerstone for establishing the error bound for Problem (2). It is worth noting that each critical point is not an isolated point but a union of connected sets (see 2). This brings significant difficulty to compute the distance $\mathrm{dist}(\bm{W},\cal W)$ from a point $\bm{W}$ to the critical point set. Such an intricacy also underscores the importance of a thorough and precise characterization of the critical point set to facilitate rigorous error-bound analysis.
@@ 2.2 Error Bound for Deep Linear Networks
In this subsection, we present our main result, which establishes the error bound for Problem (2). Before we proceed, it is important to point out that there are some degenerate cases under which the error bound does not hold. For example, when $L=2$, we consider
$\min_{\bm{W}_{1},\bm{W}_{2}\in\mathbb{R}^{d\times d}}F(\bm{W})=\frac{1}{2}\|\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\|_{F}^{2}+\frac{1}{2}\lambda\left(\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}\right).$
This, together with 1, yields that the set of critical points is $\mathcal{W}=\{(\bm{0},\bm{0})\}$. Now, consider $\bm{W}_{1}=\bm{W}_{2}=x\bm{I}_{d}$. We compute
$\nabla_{\bm{W}_{1}}F(\bm{W})=\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)+\lambda\bm{W}_{1}=x^{3}\bm{I}_{d},$
$\nabla_{\bm{W}_{2}}F(\bm{W})=\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)\bm{W}_{1}^{T}+\lambda\bm{W}_{2}=x^{3}\bm{I}_{d}.$
Therefore, we have $\|\nabla F(\bm{W})\|_{F}^{2}=2dx^{6}$. Moreover, we compute $\mathrm{dist}^{2}(\bm{W},\mathcal{W})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}=2dx^{2}$. Consequently, we have $\|\nabla F(\bm{W})\|_{F}=\mathrm{dist}^{3}(\bm{W},\mathcal{W})/(2d)$, and thus the error bound does not hold as $x\to 0$. To avoid such degenerate cases, we impose the following conditions on the relationship between the regularization parameters and the data matrix $\bm{Y}$.
@@
$\lambda_{1}\lambda_{2}\neq y_{i}^{2},\ \forall i\in[r_{Y}].$ (9)
@@
$\lambda_{1}\dots\lambda_{L}\neq y_{i}^{2(L-1)}\left(\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right)^{-2(L-1)},\ \forall i\in[r_{Y}].$ (10)
Notably, this assumption provides sufficient and necessary conditions under which the error bound holds for all critical points of Problem (2). Indeed, if 2 holds, the error bound can be rigorously established (see 2). Conversely, if 2 is violated, we can show that the error bound fails to hold (see Appendix B). This dual perspective highlights the critical role of 2 in establishing the error bound for Problem (2). Now, under Assumptions 1 and 2, we are ready to prove the error bound for Problem (2).
Theorem 2.
Suppose that Assumptions 1 and 2 hold. There exist constants $\epsilon_{1},\kappa_{1}>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W})\leq\epsilon_{1}$,
$\mathrm{dist}(\bm{W},\mathcal{W})\leq\kappa_{1}\|\nabla F(\bm{W})\|_{F}.$ (11)
This theorem is significant as it establishes the fact that the error bound holds at any critical point of Problem (2) under 2. Importantly, the constants $\kappa_{1},\epsilon_{1}$ can be explicitly derived through our proofs; see Section 3.3. Now, we discuss some implications and the related work on this result. First, the established error bound can be used to derive other regularity conditions for analyzing the convergence behavior of first-order methods. The PŁ inequality and quadratic growth of Problem (2) can be obtained in 1 when the error bound holds, whose proof is deferred to Section C.1.
@@
Suppose that $\bm{W}^{*}$ is a critical point such that (11) holds for all $\bm{W}$ satisfying $\|\bm{W}-\bm{W}^{*}\|_{F}\leq\epsilon_{1}$. The following statements hold:
(i) There exists a constant $\mu_{1}>0$ such that
@@
(ii) If $\bm{W}^{*}$ is a local minimizer, there exists a constant $\mu_{2}>0$ such that
$\mathrm{dist}^{2}(\bm{W},\mathcal{W})\leq\mu_{2}\left(F(\bm{W})-F(\bm{W}^{*})\right).$
Second, combining the error bound with other algorithm-dependent conditions, such as sufficient decrease, cost-to-go estimate, and safeguard, we obtain local linear convergence of first-order methods for solving Problem (2) (see 7). In contrast to algorithm-specific convergence rate analyses in [4, 6, 19, 32, 43, 49], which are tailored to the dynamics of individual algorithms, the error-bound-based framework provides a unified approach. It applies not only to gradient descent but also to a broad class of first-order methods capable of optimizing Problem (2). It is also worth mentioning that our experimental results in Section 5 demonstrate that GD achieves linear convergence to both optimal and non-optimal critical points across various network architectures, which further supports our theoretical findings.
Third, we emphasize our technical contributions to establishing the error bound. Unlike prior works [30, 21, 22, 42], where the critical point set $\cal W$ had a relatively simple structure that allows for explicit computation of the distance $\mathrm{dist}(\bm{W},\mathcal{W})$, the presence of orthogonal permutations and hierarchical structures in $\cal W$ (see 1) poses a significant challenge to our analysis. To address this, we carefully construct an intermediate point $\hat{\bm{W}}$ that leverages the singular vectors of $\bm{W}$ and singular values of $\bm{W}^{*}\in\cal W$. Then, we bound $\mathrm{dist}(\hat{\bm{W}},\mathcal{W})$ and $\|\bm{W}-\hat{\bm{W}}\|_{F}$ by the gradient norm at $\bm{W}$, respectively. Combining these bounds with the triangle inequality yields (11). Notably, this construction technique provides a new approach to showing the error bound of non-convex problems with a complicated solution set.

### v3

@@ 2 Main Results
To begin, we define the set of critical points of Problem (1) as follows:
$\mathcal{W}_{F}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla F(\bm{W})=\bm{0}\right\}.$
Let $\bm{Y}=\bm{U}_{Y}\bm{\Sigma}_{Y}\bm{V}_{Y}^{T}$ be an SVD of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$, where $\bm{U}_{Y}\in\mathcal{O}^{d_{L}}$ and $\bm{V}_{Y}\in\mathcal{O}^{d_{0}}$. In addition,
$\bm{\Sigma}_{Y}=\mathrm{BlkD}\left(\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}}),\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right)\in\mathbb{R}^{d_{L}\times d_{0}},$ (3)
where $d_{\min}:=\min\{d_{0},d_{L}\}$ and $y_{1}\geq y_{2}\geq\dots\geq y_{d_{\rm min}}\geq 0$ are singular values. Here, if $d_{\min}=d_{0}$ (resp., $d_{\min}=d_{L}$), the lower diagonal block vanishes, leaving a $\bm{0}$ block of the appropriate size at the bottom (resp., on the right). Specifically, if $d_{\min}=d_{0}$, $\bm{\Sigma}_{Y}=\begin{bmatrix}\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}})\\ \bm{0}_{(d_{L}-d_{\min})\times d_{0}}\end{bmatrix}$; otherwise, $\bm{\Sigma}_{Y}=\begin{bmatrix}\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}})&\bm{0}_{d_{L}\times(d_{0}-d_{\min})}\end{bmatrix}$. To handle the repeated singular values, we introduce the following setup. Let $r_{Y}:=\mathrm{rank}(\bm{Y})$ and $p_{Y}$ denote the number of distinct positive singular values of $\bm{Y}$. That is, there exist indices $s_{0},s_{1},\dots,s_{p_{Y}}$ such that $0=s_{0}<s_{1}<\dots<s_{p_{Y}}=r_{Y}$ and
$\textstyle y_{s_{0}+1}=\dots=y_{s_{1}}>y_{s_{1}+1}=\dots=y_{s_{2}}>\dots>y_{s_{p_{Y}}+1}=\cdots=y_{d_{\text{min}}}=0.$ (4)
Note that if $r_{Y}=d_{\min}$, then $y_{s_{p_{Y}}}=y_{d_{\min}}>0$. Let $h_{i}:=s_{i}-s_{i-1}$ denote the multiplicity of the $i$-th largest positive singular value for each $i\in[p_{Y}]$. Throughout the paper, we consistently use the above notation.
2.1 Characterization of the Critical Point Set
To begin, we introduce the following assumption on the widths of network layers.
@@ 2.1 Characterization of the Critical Point Set
The network widths satisfy $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$.
This assumption ensures that the width of each hidden layer is no less than that of the input or output layer. Notably, it aligns with common practices in deep learning, where hidden layers are typically designed to be wide to have sufficient capacity to capture complex data representations [16, 35]. In addition, this assumption is widely used in the literature to analyze the convergence behavior of GD for optimizing deep networks; see, e.g., [1, 11, 12, 26]. Then we derive the closed-form expression of the critical point set $\mathcal{W}_{F}$ as follows:
Theorem 1 (Characterization of critical points).
Suppose that Assumption 1 holds. It holds that $\bm{W}\in\mathcal{W}_{F}$ if and only if
$\bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right)\bm{V}_{Y}^{T},$ (5a)
@@
$\bm{W}_{L}=\bm{U}_{Y}\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},$ (5c)
where $\bm{Q}_{l}\in\mathcal{O}^{d_{l-1}}$ for all $l=2,\dots,L$, $\bm{O}_{i}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$, $\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}}$, $\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}$, and $\bm{\Sigma}_{l}=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma})/\sqrt{\lambda_{l}},\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for each $l\in[L]$ with $\bm{\sigma}\in\mathbb{R}^{r_{Y}}$ satisfying
$\textstyle\sigma_{i}^{2L-1}-\sqrt{\lambda}y_{i}\sigma_{i}^{L-1}+\lambda\sigma_{i}=0,\ \sigma_{i}\geq 0,\ \forall i\in[r_{Y}],\ \text{where}\ \lambda:=\prod_{l=1}^{L}\lambda_{l}.$ (6)
We remark that this theorem is a special case of [6, Theorem 2.1], which provides a closed-form characterization of the critical point set of Problem (1) without any conditions on the network widths. In contrast, by leveraging Assumption 1, we obtain a simplified and more explicit description of the critical point set in this theorem. This characterization serves as a cornerstone for establishing the error bound for Problem (1). In general, these critical points are not isolated points but form a union of connected sets (see 1). This brings significant difficulty in computing the distance $\mathrm{dist}(\bm{W},\mathcal{W}_{F})$ from a point $\bm{W}$ to the critical point set, as this distance does not admit a closed-form solution due to the complex form of $\mathcal{W}_{F}$. One of our key contributions is to develop a method that overcomes this difficulty.
@@ 2.2 Error Bound for Deep Linear Networks
Before we proceed, it is worth noting that there are some degenerate cases under which the error bound does not hold. For example, when $L=2$, we study
$\min_{\bm{W}_{1},\bm{W}_{2}\in\mathbb{R}^{d\times d}}F(\bm{W})=\|\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\|_{F}^{2}+\lambda\left(\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}\right).$
Applying 1 to the above problem yields $\mathcal{W}_{F}=\{(\bm{0},\bm{0})\}$. Now, consider $\bm{W}_{1}=\bm{W}_{2}=x\bm{I}_{d}$. We compute
$\nabla_{1}F(\bm{W})=2\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)+2\lambda\bm{W}_{1}=2x^{3}\bm{I}_{d},$
$\nabla_{2}F(\bm{W})=2\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)\bm{W}_{1}^{T}+2\lambda\bm{W}_{2}=2x^{3}\bm{I}_{d}.$
Therefore, we have $\|\nabla F(\bm{W})\|_{F}^{2}=8dx^{6}$. Moreover, we compute $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}=2dx^{2}$. Consequently, we have $\sqrt{d}\|\nabla F(\bm{W})\|_{F}=\mathrm{dist}^{3}(\bm{W},\mathcal{W})$. Obviously, the error bound (2) does not hold as $x\to 0$. To avoid such degenerate cases, we impose the following conditions on the relationship between the regularization parameters and the data matrix $\bm{Y}$.
@@
$\lambda:=\lambda_{1}\lambda_{2}\neq y_{i}^{2},\ \forall i\in[r_{Y}].$ (7)
@@
$\lambda:=\lambda_{1}\cdots\lambda_{L}\neq y_{i}^{2(L-1)}\left(\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right)^{-2(L-1)},\ \forall i\in[r_{Y}].$ (8)
Notably, this assumption provides necessary and sufficient conditions under which the error bound holds for all critical points of Problem (1). Specifically, if Assumption 2 does not hold, the error bound fails to hold (see Appendix A). Conversely, if Assumption 2 holds, the error bound holds as follows:
Theorem 2 (Error bound).
Suppose that Assumptions 1 and 2 hold. There exist constants $\epsilon_{1},\kappa_{1}>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon_{1}$, we have
$\mathrm{dist}\left(\bm{W},\mathcal{W}_{F}\right)\leq\kappa_{1}\|\nabla F(\bm{W})\|_{F}.$ (9)
This theorem is significant as it shows that the error bound holds at any critical point of Problem (1) under Assumptions 1 and 2. Importantly, the constants $\kappa_{1},\epsilon_{1}$ can be explicitly derived through our proofs (see 2, 3, and (125) and (126) in the proof of 2). Now, we discuss some implications of our result. First, the established error bound can be used to derive other regularity conditions, such as PŁ inequality (see (10)) and quadratic growth (see (11)), for analyzing the convergence behavior of first-order methods. This result is well known; see, e.g., [36].
@@
Suppose that $\bm{W}^{*}$ is a critical point and (9) holds for all $\bm{W}$ satisfying $\|\bm{W}-\bm{W}^{*}\|_{F}\leq\epsilon_{1}$.
(PŁ inequality) There exists a constant $\mu_{1}>0$ such that
@@
(ii) (Quadratic growth) If $\bm{W}^{*}$ is a local minimizer, there exists a constant $\mu_{2}>0$ such that
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})\leq\mu_{2}\left(F(\bm{W})-F(\bm{W}^{*})\right).$ (11)
Second, by combining the error bound with other algorithm-dependent conditions, such as sufficient decrease, cost-to-go estimate, and safeguard, we can establish local linear convergence of first-order methods for solving Problem (1) (see Appendix C). In contrast to algorithm-specific convergence rate analyses in [4, 7, 20, 32, 44], the error-bound-based framework provides a unified analytical approach. In particular, it applies not only to GD but also to a broad class of first-order methods capable of optimizing Problem (1). Moreover, our experimental results in Section 4 demonstrate that GD achieves linear convergence to critical points, which further supports our theoretical findings.
Third, we emphasize our technical contributions to establishing the error bound. Unlike prior works [30, 21, 22, 43], where the critical point set $\mathcal{W}_{F}$ has a relatively simple structure that allows for closed-form computation of the distance $\mathrm{dist}(\bm{W},\mathcal{W}_{F})$, the presence of orthogonal permutations and hierarchical structures in $\mathcal{W}_{F}$ (see 1) prevents such computation. To address this challenge, we develop a suite of new techniques (see Section 3.2.2), which offer a new framework for establishing error bounds of non-convex problems with complex solution sets.